1/27=9to the 5x-5 power
Answer by stanbon(75887)

1/27=9to the 5x-5 power
---
3^-3 = (3^2)^(5x-5)
3^-3 = 3^(15x-15)
---
15x-15 = -3
15x = 12
x = 4/5
=====================
Cheers,
Stan H.